OpenStudy (anonymous):

Hey guys, having trouble grasping logs at the moment...very basic stuff! Just wondering how you'd use a calculator to find * when 10*=40?

OpenStudy (anonymous):

*= 4 since, \( 10 \times 4 = 40 \)

OpenStudy (anonymous):

sorry its 10 to the power of 4...I couldnt find a symbol to make it clear on the computer!

OpenStudy (anonymous):

Alight so, you want to solve for \( x \) in \( 10^x=4 \), isn't ?

OpenStudy (anonymous):

exactly!

OpenStudy (anonymous):

Okay so, \( 10^x =4 \implies x = \log_{10} 4 \approx 0.602 \) Here is a cool website that you would use: http://www.wolframalpha.com/input/?i=Log%5B10%2C4%5D

OpenStudy (anonymous):

Thats great, thanks so much!! :)

OpenStudy (anonymous):

Glad to help :-)

OpenStudy (anonymous):

hi every body i waant to know that what is active and passive elements in elctronics circuits?